# MuseGraph: Graph-oriented Instruction Tuning of Large Language Models for Generic Graph Mining

### 02 Mar 2024 v1

@@ 1. Introduction
Recently, Graph Neural Networks (GNNs) have become widely used in modeling attributed graphs (hamilton2017inductive, ; wang2019heterogeneous, ; cui2020adaptive, ). However, they need to be re-trained whenever applied to different graph datasets and tasks. Inspired by the great success of Large Language Models (LLMs), the combination of GNNs and LLMs aims to enhance the processing of text-attributed graphs, which can be categorized into two main approaches. The first line tries to train GNNs with LLM-enhanced features (e.g., OFA (liu2023one, ), ALL-in-One (sun2023all, ), GIMLET (zhao2023gimlet, ), ExpAsFeat (he2023explanations, ), LLMForGraph (chen2023exploring, ) and PRODIGY (huang2023prodigy, )) The second line tunes LLMs for graph applications (e.g., InstructGLM (ye2023natural, ), NLGraph (wang2024language, ) and LMMoL (qian2023can, ), GPT4GRAPH (guo2023gpt4graph, )).
@@
Challenge II: How to automatically generate diverse instructions? Creating a diverse set of high-quality instructions for fine-tuning LLMs is essential for generic graph mining tasks. However, these instructions are often neither readily available nor economically feasible to produce manually. While advanced language models like GPT-4 possess the capability to generate diverse instructions, it remains unknown how to effectively leverage these generative abilities to produce relevant, task-specific instructions.
@@
To tackle these challenges, we propose Graph-oriented Instruction Tuning of Large Language Models for Generic Graph Mining (MuseGraph), which consists of three pivotal steps: (i) Development of Compact Graph Descriptions, where we introduce a novel “node energy” metric to texualize graphs with essential semantic and structural details under limited tokens; (ii) Generation of Diverse Instructions, which invokes the advanced generative abilities of GPT-4 to create CoT-based instruction packages tailored for various graph tasks, thus enriching LLMs’ capabilities in understanding and analyzing graph data without the expense of manual instruction crafting; (iii) Graph-aware Instruction Tuning, which introduces a dynamic instruction package allocation strategy based on the specific needs of each graph mining task, ensuring comprehensive and effective LLM tuning.
@@
  • Formulation of generic graph mining. We establish a generic graph framework that effectively transforms graph structures into LLM-friendly formats while preserving the generative capabilities necessary for diverse graph-oriented tasks
  • Effective Model Designs. We design and implement a set of models and mechanisms including the development of compact graph descriptions, automatically generating diverse, task-specific instructions, and graph-aware instruction tuning, targeting a generic graph model across tasks and datasets.
@@ 2.1. Semantic-rich Graph Representation Learning
Graph Representation Learning has emerged as a key technique for the complex structures of networks with abundant attributes (DBLP:journals/debu/HamiltonYL17, ; DBLP:journals/tkde/CaiZC18, ; DBLP:journals/tnn/WuPCLZY21, ; zhou2020graph, ; DBLP:journals/tbd/WangBSFYY23, ). Among existing node embedding methods, many had analyzed and utilized the great promise of random walks (RWs) in capturing the topological structures of graphs (perozzi2014deepwalk, ; grover2016node2vec, ; dong2017metapath2vec, ; ivanov2018anonymous, ). However, the above methods ignored the abundant attribute information surrounding the nodes and edges (mao2023hinormer, ). Recently, Graph neural networks (GNNs) for learning node representations through aggregating information from neighboring nodes on graphs (kipfsemi, ; hamilton2017inductive, ; yang2020relation, ). However, most existing GNNs were established in a supervised learning setting, which requires abundant task-specific labeled data that may not be available in real-world applications (zhu2021graph, ; cooray2022graph, ), and the embeddings they learned are not generalizable across different downstream tasks (zhang2018link, ), which had to be re-trained whenever applied to different graph tasks and dataset.
@@ 2.2. Leveraging LLMs for Graph Mining
NLGraph (wang2024language, ) $\sim$
GPT4GRAPH (guo2023gpt4graph, ) $\sim$
GraphGPT (guo2023gpt4graph, ) $\sim$
InstructGLM (ye2023natural, ) $\sim$
@@
Motivated by the impressive strides in Large Language Models (LLMs), combining Graph Neural Networks (GNNs) and LLMs is creating substantial progress in handling complex text-attributed graphs (li2024survey, ; chen2023label, ; chen2023exploring, ). Methods leveraging this combination fall into two broad categories (shown in Table 1).
@@
These models primarily utilized LLMs to process textual content within graphs and tend to be tailored to specific domains, often failed to be applied for cross-task and cross-dataset applications and lost the generation capability from LLMs. Although these models effectively focus on leveraging LLMs’ textual prowess, most of them are domain-specific graph models, which fail to perform cross-task and cross-dataset.
The second category denoted as LLM* in Table 1, mainly focuses on tuning LLMs for graph-related tasks. For example, InstructGLM (ye2023natural, ), NLGraph (wang2024language, ), and GPT4GRAPH (guo2023gpt4graph, ) leveraged LLMs to understand and analyze graph data through textual descriptions. GraphGPT (tang2023graphgpt, ) with one checkmark on GNN* and a partial checkmark on LLM* jointly trained task-specific GNN and one projector on the top of LLMs, which performed well across different datasets. Note that, the partial checkmark for some methods means that they can be extended to cross-dataset targets, but this is not their main research focus and they have no design mechanism to achieve it. Compared with GNN*-based methods, LLM*-based methods have the potential to perform generic graph mining with powerful cross-tasks and cross-datasets. However, such power remains largely under-explored. Table 1 summarizes the characteristics of different models. We can see that our MuseGraph is the most generic model among existing works.
@@ 3.1. Overview of Our Framework
Objective: In this paper, we aim to pursue an ideal graph that can fascinate a more effective and generic approach via graph-oriented instruction tuning of LLMs, through which we wish to seamlessly integrate the strengths of GNNs and LLMs for graph mining across different tasks and dataset
Overview: To achieve this goal, we propose MuseGraph framework, which comprises three major components. Firstly, we develop a compact graph description mechanism that captures critical semantic and structural details within the constraints of language tokens. Secondly, we generate a diverse range of instructions through the reasoning capabilities of LLMs, thus facilitating task-specific, Chain-of-Thought-based instruction packages for various graph tasks. Thirdly, we adopt a graph-aware instruction tuning, utilizing a dynamic allocation strategy for instruction packages tailored to the unique demands of each graph task. The overall model architecture is shown in Figure 2, and we elaborate on the main components.
@@ 3.2. Compact Graph Description
Leveraging the capabilities of Large Language Models (LLMs) for graphs presents a unique set of challenges, primarily due to LLMs’ inherent limitations in directly inputting graph structures under the token limitation. Therefore, it is non-trivial to automatically extract key information from graphs under the token limitations, which require a compact description including complex node and edge attributes along with structural details.
@@
Given LLM token limits and the varied contribution of neighbors and walks to node understanding, we further develop an adaptive input generation mechanism to ensure the compactness of the description. Specifically, we first design a “node energy” metric, $H(v_{i})$ assessing node information from two perspectives: token count in attributes and node degree count. This metric enables us to effectively filter and select neighbor nodes and walk nodes, prioritizing those that are abundant in semantic information and possess a significant number of neighboring nodes, thus enhancing the expressiveness of the graph description. The calculation of node energy $H(v_{i})$ is formulated as follows:
@@
where $v_{i}$ is the target node, ${T(v_{i})}$ is the number of node tokens processed by a tokenizer and ${D(v_{i})}$ is the number of node degrees.
10:  Attributed graph $\mathcal{G}$ with $N$ node, token count set $\mathcal{T}$,
@@
Based on $H(v_{i})$, our method strategically incorporates neighbors and walks, as depicted in Algorithm 1. A neighbor, $v_{i}$, is chosen to describe the target $*v*$ if its $H(v_{i})$ surpasses $H(v^{*})$, ensuring the included node can provide supplemental information. The process of including walks concludes when encountering a node whose $H(v_{i})$ does not meet the threshold set by the target’s $H(v^{*})$, thus refining the input to maximize relevance within the constraints of the token limit. To further demonstrate the adaptive input generation for neighbors and walks tailored to each node, we present an example in Figure 3). Node $v_{1}$’s with a relatively low $H(v_{1})$ value, includes a wide range of neighbors to capture more context, excluding $v_{4}$ due to its even lower $H(v_{4})$. Given the token limitations, only two walks are sampled for $v_{1}$ to maintain a compact input. Conversely, $v_{2}$ with its higher $H(v_{2})$, inherently carries more information, prompting the selection of fewer neighbors. $v_{1}$ is excluded since $H(v_{1})<H(v_{2})$. This frees up tokens to detail more walks for $v_{2}$. The $H$ metric thus finely balances neighbor and walk inclusion for each node, marrying information-rich and token-efficient. The process culminates in the textualization of each node’s key information, producing a tailored and compact graph description as depicted in the upper right of Figure 2. Note that, a graph-related task can involve multiple nodes. In this case, we further allocate each node to satisfy the token limit requirement by calculating the ratio of all node energies using the softmax function.
@@ 3.3. Diverse Instruction Generation
Specifically, we first design diverse CoT templates for targeted popular graph tasks, such as node classification, link prediction, and graph-to-text (shown in A Table 7). Then, we prompt GPT-4 to generate a small number of CoT-based instructions across diverse tasks. This approach distills GPT-4’s vast knowledge base to augment the reasoning and analytical abilities of our MuseGraph, as depicted in Figure 4.
@@ 3.4. Graph-aware Instruction Tuning
To achieve the balance between effectiveness and efficiency, we adopt a graph-aware instruction tuning mechanism, which can sufficiently utilize diverse and high-quality instructions for further fine-tuning LLMs (ouyang2022training, ). Specifically, we adopt a general LLM LLaMA-V1-7B (abbr. LLaMA) (touvron2023llama, ) with LoRA (hu2021lora, ) as our starting point for fine-tuning. Then, based on diverse instruction set =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{N}\}$ as model input, we adopt the negative log-likelihood loss as the fine-tuning objective (achiam2023gpt, ; ye2023natural, ), which is formulated as follows:
@@ 4.2. Overall Performance Comparison (RQ1)
To evaluate our MuseGraph framework, we conducted extensive testing across node classification and graph-to-text tasks using five real-world datasets.
@@
In the node classification task, MuseGraph demonstrates significant performance gains, ranging from reasonably large (12.33% over GAT on MIMIC-III) to significantly large (19.67% over GLNN on Cora). In the graph-to-text task, the performance gains of MuseGraph over all GNN-based baselines range from 1.02% (achieved in METEOR on WebNLG) to 57.89% (achieved in BLEU-4 on AGENDA). Note that, GNN-based methods (e.g., GLNN and GAT) only achieve second-runner in the node classification task. Since the generative and reasoning abilities of GNN-based methods are weaker than LLM-based ones in generative tasks, they can not achieve the second-runner in the graph-to-text task. In particular, GLNN is optimized for attribute-rich datasets (e.g., Arxiv and Cora) through its knowledge distillation mechanisms, while GAT is tailored for datasets with basic attributes (e.g., MIMIC-III) through its ability to weigh nodes differently. Compared with GLNN and GAT, our MuseGraph with graph-ware instruction tuning with the diverse CoT-based instruction packages can effectively comprehend graph data and achieve accurate predictions of various downstream tasks.
@@
Note that, MuseGraph can sometimes exceed all baselines with ROUGE-L metric, achieving an average performance gain of 22.40% over the second-runner T5 on average. By extracting the compact graph description by adaptively allocation the neighbors and walks under the input token limitation, MuseGraph adeptly extracts a compact graph description. This enables it to accurately grasp both the semantic and structural nuances of graphs through a graph-aware fine-tuning mechanism. Given that ROUGE-L prioritizes the longest common subsequence between generated and reference texts, MuseGraph’s focus on key graph information translates into texts that are both accurate and coherent. This capability is particularly valued in ROUGE-L assessments, which do not require strict adherence to the sequence of words and phrases at the word or phrase level. As a result, MuseGraph’s generated texts can achieve high scores even when the structure is more flexible.
@@ 4.3. Ablation Studies (RQ2)
To study the effectiveness of different allocations of CoT-based instruction packages across various tasks and datasets, we conduct ablation studies for node classification on Arxiv, which is based on our proposed compact graph description for inputting graph. Using LLaMA as our baseline model, the studies evaluated our MuseGraph across various configurations, while maintaining an equal total count of instruction packages. We study our MuseGraph as follows:
  • 1 Task+1 Dateset: Utilizes our proposed CoT-based instructions for node classification on Arxiv.
  • 2 Task+1 Dateset: Incorporates instructions from both node classification and link prediction on Arxiv.
  • 2 Task+2 Dateset: Expands to include node classification and link prediction instructions from both Arxiv and Cora.
  • 3 Task+3 Dataset: Engages with instructions across node classification, link prediction on Arxiv and Cora, and graph-to-text on AGENDA..
@@
With adding more types of tasks and datasets, our proposed MuseGraph can achieve further improvements, where 3 Task+3 Dataset overperforms 2 Task+2 Dataset ranges from 3.03% in Macro-F1 to 3.94% in Weighted-F1. Such observation strongly indicates that the dynamic instruction package allocation strategy can enhance the accuracy of graph tasks and avert catastrophic forgetting, where LLMs might regress in their generative abilities.

### 13 Mar 2024 v2

@@ 1. Introduction
Recently, Graph Neural Networks (GNNs) have become widely used in modeling attributed graphs (hamilton2017inductive, ; wang2019heterogeneous, ; cui2020adaptive, ). However, they need to be re-trained whenever applied to different graph tasks and datasets. Inspired by the great success of Large Language Models (LLMs), the combination of GNNs and LLMs aims to enhance the processing of text-attributed graphs, which can be categorized into two main approaches. The first line tries to train GNNs with LLM-enhanced features (e.g., OFA (liu2023one, ), ALL-in-One (sun2023all, ), GIMLET (zhao2023gimlet, ), ExpAsFeat (he2023explanations, ), LLMForGraph (chen2023exploring, ) and PRODIGY (huang2023prodigy, )). The second line tunes LLMs for graph applications (e.g., InstructGLM (ye2023natural, ), NLGraph (wang2024language, ), LMMoL (qian2023can, ), and GPT4GRAPH (guo2023gpt4graph, )).
@@
Challenge II: How to automatically generate diverse instructions? Creating a diverse set of high-quality instructions for fine-tuning LLMs is essential for generic graph mining tasks. However, these instructions are often neither readily available nor economically feasible to produce manually. While advanced language models like GPT-4 possess the capability to generate diverse instructions, it remains unknown how to effectively leverage these generative abilities to produce relevant and task-specific instructions.
@@
To tackle these challenges, we propose Graph-oriented Instruction Tuning of Large Language Models for Generic Graph Mining (MuseGraph), which consists of three pivotal steps: (i) Development of Compact Graph Descriptions, where we introduce a novel “node energy” metric to textualize graphs with essential semantic and structural details under limited tokens; (ii) Generation of Diverse Instructions, which invokes the advanced generative abilities of GPT-4 to create CoT-based instruction packages tailored for various graph tasks, thus enriching LLMs’ capabilities in understanding and analyzing graph data without the expense of manual instruction crafting; (iii) Graph-aware Instruction Tuning, which introduces a dynamic instruction package allocation strategy based on the specific needs of each graph mining task, ensuring comprehensive and effective LLM tuning.
@@
  • Formulation of generic graph mining. We establish a generic graph framework that effectively transforms graph structures into LLM-friendly formats while preserving the generative capabilities of necessary for diverse graph-oriented tasks.
  • Effective Model Designs. We design and implement a set of models and mechanisms including the development of compact graph descriptions, automatically generating diverse, task-specific instructions, and graph-aware instruction tuning, targeting a generic graph model across different tasks and datasets.
@@ 2.1. Semantic-rich Graph Representation Learning
Graph representation learning has emerged as a key technique for the complex structures of networks with abundant attributes (DBLP:journals/debu/HamiltonYL17, ; DBLP:journals/tkde/CaiZC18, ; DBLP:journals/tnn/WuPCLZY21, ; zhou2020graph, ; DBLP:journals/tbd/WangBSFYY23, ). Among existing node embedding methods, many have analyzed and utilized the great promise of random walks (RWs) in capturing the topological structures of graphs (perozzi2014deepwalk, ; grover2016node2vec, ; dong2017metapath2vec, ; ivanov2018anonymous, ). However, the above methods ignore the abundant attribute information surrounding the nodes and edges (mao2023hinormer, ). Recently, Graph neural networks (GNNs) learn node representations through aggregating information from neighboring nodes on graphs (kipfsemi, ; hamilton2017inductive, ; yang2020relation, ). However, most existing GNNs are established in a supervised learning setting, which requires abundant task-specific labeled data that may not be available in real-world applications (zhu2021graph, ; cooray2022graph, ), and the embeddings that they learned are not generalizable across different downstream tasks (zhang2018link, ), which has to be re-trained whenever applied to different graph tasks and datasets.
@@ 2.2. Leveraging LLMs for Graph Mining
NLGraph (wang2024language, )
$\sim$
GPT4GRAPH (guo2023gpt4graph, )
$\sim$
GraphGPT (guo2023gpt4graph, )
$\sim$
InstructGLM (ye2023natural, )
$\sim$
@@
Figure 2. The overall framework of MuseGraph.
Motivated by the impressive strides in Large Language Models (LLMs), combining Graph Neural Networks (GNNs) and LLMs is creating substantial progress in handling complex text-attributed graphs (li2024survey, ; chen2023label, ; chen2023exploring, ). Existing methods leveraging this combination mainly fall into two broad categories (shown in Table 1).
@@
These models primarily utilized LLMs to process textual content within graphs and tend to be tailored to specific domains, while they often failed to be applied for cross-task and cross-dataset applications and lost the generation capability from LLMs. Although these models effectively focus on leveraging LLMs’ textual prowess, most of them are domain-specific graph models, which fail to perform cross-task and cross-dataset.
The second category denoted as LLM* in Table 1, mainly focuses on tuning LLMs for graph-related tasks. For example, InstructGLM (ye2023natural, ), NLGraph (wang2024language, ), and GPT4GRAPH (guo2023gpt4graph, ) leveraged LLMs to understand and analyze graph data through textual descriptions. GraphGPT (tang2023graphgpt, ) with one checkmark on GNN* and a partial checkmark on LLM* jointly trained task-specific GNN and one projector on the top of LLMs, which performed well across different datasets. Note that, the partial checkmark $\sim$  for some methods means that they can be extended to cross-dataset targets, but this is not their main research focus and they have no design mechanism to achieve it. Compared with GNN*-based methods, LLM*-based methods have the potential to perform generic graph mining with powerful cross-task and cross-dataset. However, such power remains largely under-explored. Table 1 summarizes the characteristics of different models. We can see that our MuseGraph is the most generic model among existing works.
@@ 3.1. Overview of Our Framework
Objective: In this paper, we aim to pursue an ideal graph that can seamlessly integrate the strengths of GNNs and LLMs via graph-oriented instruction tuning of LLMs, through which we wish to facilitate a more effective and generic approach for graph mining across different tasks and datasets.
Overview: To achieve this goal, we propose MuseGraph framework, which comprises three major components. Firstly, we develop a compact graph description mechanism that captures critical semantic and structural details within the constraints of language token limitations. Secondly, we generate a diverse range of instructions through the reasoning capabilities of LLMs, thus facilitating task-specific and Chain-of-Thought-based instruction packages for various graph tasks. Thirdly, we adopt a graph-aware instruction tuning, utilizing a dynamic allocation strategy for instruction packages tailored to the unique demands of each graph task. The overall model architecture is shown in Figure 2, and we elaborate on the three main components.
@@ 3.2. Compact Graph Description
Leveraging the capabilities of Large Language Models (LLMs) for graphs presents a unique set of challenges, primarily due to LLMs’ inherent limitations in directly inputting graph structures under the token limitation. Therefore, it is non-trivial to automatically extract key information from graphs under the token limitations, which requires a compact description including complex node and edge attributes along with structural details.
@@
Given LLM token limitations and the varied contribution of neighbors and walks to node understanding, we further develop an adaptive input generation mechanism to ensure the compactness of the description. Specifically, we first design a “node energy” metric, $H(v_{i})$ assessing node information from two perspectives: token count in node attributes and node degree count. This metric enables us to effectively filter and select neighbor nodes and walk nodes, prioritizing those that are abundant in semantic information and possess a significant number of neighboring nodes, thus enhancing the expressiveness of the graph description. The calculation of node energy $H(v_{i})$ is formulated as follows:
@@
where $v_{i}$ is the target node, ${T(v_{i})}$ is the number of node tokens processed by a tokenizer, and ${D(v_{i})}$ is the number of node degrees.
10:  Attributed graph $\mathcal{G}$ with $N$ nodes, token count set $\mathcal{T}$,
@@
Based on $H(v_{i})$, our method strategically incorporates neighbors and walks, as depicted in Algorithm 1. A neighbor $v_{i}$ is chosen to describe the target $v^{*}$ if its $H(v_{i})$ surpasses $H(v^{*})$, ensuring the included node can provide supplemental information. The process of including walks concludes when encountering a node whose $H(v_{i})$ does not meet the threshold set by the target’s $H(v^{*})$, thus refining the input to maximize relevance within the constraints of the token limit. To further demonstrate the adaptive input generation for neighbors and walks tailored to each node, we present an example in Figure 3. Node $v_{1}$’s with a relatively low $H(v_{1})$ value, includes a wide range of neighbors to capture more context, excluding $v_{4}$ due to its even lower $H(v_{4})$. Given the token limitations, only two walks are sampled for $v_{1}$ to maintain a compact input. Conversely, $v_{2}$ with its higher $H(v_{2})$, inherently carries more information, prompting the selection of fewer neighbors. $v_{1}$ is excluded since $H(v_{1})<H(v_{2})$. This frees up tokens to detail more walks for $v_{2}$. The $H$ metric thus finely balances neighbor and walk inclusion for each node, marrying information-rich and token-efficient characteristics. The process culminates in the textualization of each node’s key information, producing a tailored and compact graph description as depicted in the upper right of Figure 2. Note that, a graph-related task can involve multiple nodes. In this case, we further allocate each node to satisfy the token limit requirement by calculating the ratio of all node energies using the softmax function.
@@ 3.3. Diverse Instruction Generation
Specifically, we first design diverse CoT templates for targeted popular graph tasks, such as node classification, link prediction, and graph-to-text (shown in Appendix A.3 Table 7). Then, we prompt GPT-4 to generate a small number of CoT-based instructions across diverse tasks. This approach distills GPT-4’s vast knowledge base to augment the reasoning and analytical abilities of our MuseGraph, as depicted in Figure 4.
@@ 3.4. Graph-aware Instruction Tuning
To achieve the balance between effectiveness and efficiency, we adopt a graph-aware instruction tuning mechanism, which can sufficiently utilize diverse and high-quality instructions for further fine-tuning LLMs (ouyang2022training, ). Specifically, we adopt a general LLM LLaMA-V1-7B (abbr. LLaMA) (touvron2023llama, ) with LoRA (hu2021lora, ) as our starting point for fine-tuning. Then, based on diverse instruction set =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{D}\}$ as model input, we adopt the negative log-likelihood loss as the fine-tuning objective (achiam2023gpt, ; ye2023natural, ), which is formulated as follows:
@@ 4.2. Overall Performance Comparison (RQ1)
To evaluate our MuseGraph framework, we conduct extensive testing across node classification and graph-to-text tasks using five real-world datasets.
@@
In the node classification task, MuseGraph demonstrates significant performance gains, ranging from reasonably large (12.33% over GAT on MIMIC-III) to significantly large (19.67% over GLNN on Cora). In the graph-to-text task, the performance gains of MuseGraph over all GNN-based baselines range from 1.02% (achieved in METEOR on WebNLG) to 57.89% (achieved in BLEU-4 on AGENDA). Note that, GNN-based methods (e.g., GLNN and GAT) only achieve second-runner in the node classification task. Since the generative and reasoning abilities of GNN-based methods are weaker than LLM-based ones in generative tasks, they can not achieve the second-runner in the graph-to-text task. In particular, GLNN is optimized for attribute-rich datasets (e.g., Arxiv and Cora) through its knowledge distillation mechanisms, while GAT is tailored for datasets with basic attributes (e.g., MIMIC-III) through its ability to weigh nodes differently. Compared with GLNN and GAT, our MuseGraph with graph-aware instruction tuning with the diverse CoT-based instruction packages can effectively comprehend graph data and achieve accurate predictions of various downstream tasks.
@@
Note that, MuseGraph can sometimes exceed all baselines with ROUGE-L metric, achieving an average performance gain of 1.58% over the second-runner T5 on average. By adaptively allocating the neighbors and walks under the input token limitation, MuseGraph adeptly extracts a compact graph description. This enables it to accurately grasp both the semantic and structural nuances of graphs through a graph-aware fine-tuning mechanism. Given that ROUGE-L prioritizes the longest common subsequence between generated and reference texts, MuseGraph’s focus on key graph information translates into texts that are both accurate and coherent. This capability is particularly valued in ROUGE-L assessments, which do not require strict adherence to the sequence of words and phrases at the word or phrase level. As a result, MuseGraph’s generated texts can achieve high scores even when the structure is more flexible.
@@ 4.3. Ablation Studies (RQ2)
To study the effectiveness of different allocations of CoT-based instruction packages across various tasks and datasets, we conduct ablation studies for node classification on Arxiv, which is based on our proposed compact graph description for inputting graph. Using LLaMA as our baseline model, the studies evaluate our MuseGraph across various configurations, while maintaining an equal total count of instruction packages. We study our MuseGraph as follows:
  • 1 Task+1 Dataset: Utilizes our proposed CoT-based instructions for node classification on Arxiv.
  • 2 Task+1 Dataset: Incorporates instructions from both node classification and link prediction on Arxiv.
  • 2 Task+2 Dataset: Expands to include node classification and link prediction instructions from both Arxiv and Cora.
  • 3 Task+3 Dataset: Engages with instructions across node classification, link prediction on Arxiv and Cora, and graph-to-text on AGENDA.
@@
With adding more types of tasks and datasets, our proposed MuseGraph can achieve further improvements, where 3 Task+3 Dataset outperforms 2 Task+2 Dataset ranging from 3.03% in Macro-F1 to 3.94% in Weighted-F1. Such observation strongly indicates that the dynamic instruction package allocation strategy can enhance the accuracy of graph tasks and avert catastrophic forgetting, where LLMs might regress in their generative abilities.